OpenStudy (anonymous):

this problem is under linear equations, 2 square root of x+3 = 7. Need help?

Parth (parthkohli):

Try squaring both sides.

Parth (parthkohli):

\( \color{Black}{\Rightarrow \sqrt{x + 3}^2 = 7^2}\)

Parth (parthkohli):

You will get \(x + 3 = 49\), now subtract 3 from both sides.

OpenStudy (lgbasallote):

\[2\sqrt{x+3} = 7\] square both sides \[(2\sqrt{x+3})^2 = 7^2\] \[4(x+3) = 49\] can you do it now? or do you need more help?

Parth (parthkohli):

Oops.

Parth (parthkohli):

I missed the 2.

Parth (parthkohli):

Crappy explanation, but you get the point.

OpenStudy (anonymous):

I get it now, thank u.

OpenStudy (lgbasallote):

welcome welcome :D

Parth (parthkohli):

You're welcome!